# Paradigms of AI Evaluation: Mapping Goals, Methodologies and Culture

### v1

@@ 1 Introduction
Figure 1: The graph shows a UMAP projection of the Jaccard distance matrix, generated using the dimensions from our framework. Each point represents a surveyed paper, whose colour indicates the paradigms it belongs to. We find clusters of papers corresponding to different paradigms. An interactive version of the plot is available here.
@@
Beyond scientific research and industry practices, underappreciating the breadth of AI evaluation may impact policy, as various policy initiatives ubiquitously rely on “evaluation”. For instance, the EU AI Act states that developers of a general-purpose AI model with “high impact capabilities evaluated on the basis of appropriate technical tools and methodologies” (European Union, 2024, Art. 51(1)) are subject to additional requirements, such as “perform[ing] model evaluation […] with a view to identifying and mitigating systemic risks” (European Union, 2024, Art. 55(1)); thus, “evaluation” refers, in the same document, both to measurement of capabilities and identification of risks. Ongoing efforts [^1] Such as the Code of Practice for general-purpose AI models, https://digital-strategy.ec.europa.eu/en/policies/ai-code-practice. identifying concrete evaluation tools to satisfy those requirements should consider the full spectrum of AI evaluation practices and understand their terminological and methodological differences, to select the appropriate set of techniques for distinct regulatory purposes.
@@ 1.2 Previous Analyses of AI Evaluation
Several recent surveys have explored AI evaluation from specific perspectives. Many of these are focused on LLMs, overviewing benchmarking (Chang et al., 2024) or red-teaming (Lin et al., 2025), studying how ethical aspects are evaluated (Lyu and Du, 2025) or analysing the literature under a verification and validation perspective (Huang et al., 2024). Other surveys include Rauh et al. (2024), which focuses on the safety of generative AI systems, and Göllner and Tropmann-Frick (2023), which considers responsible AI principles. In contrast, our work focuses on a wide variety of AI systems and evaluation approaches.
Other work discussed the quality of evaluation practice and instruments (Hernández-Orallo, 2020; Burden, 2024), for instance touching on reproducibility (Burnell et al., 2023), statistical rigour (Gorman and Bedrick, 2019), validity (Subramonian et al., 2023) and representativeness and fairness (Bergman et al., 2023; Göllner and Tropmann-Frick, 2023). We do not focus on these issues here, as they apply broadly to all AI evaluation tools.
Bieger et al. (2016) proposed a normative framework for evaluating adaptive general-purpose AI, outlining the purposes of evaluation, the properties to be measured, and the challenges involved. In contrast, we take a descriptive approach to AI evaluation, mapping the landscape by identifying the advantages and limitations of different methodologies without prescribing any particular approach. More recently, Cohn et al. (2022) defined different “facets” of evaluation instruments. While there is some overlap with our dimensions of analysis, several of their facets refer to validity and consistency, properties that are broadly relevant across all paradigms and are hence excluded from our analysis. While Cohn et al. (2022) apply their framework to 23 evaluation works, neither Bieger et al. (2016) nor Cohn et al. (2022) identify distinct paradigms of evaluation or characterise their defining features. Further, these works do not survey the larger and rapidly evolving AI evaluation landscape shaped by recent advancements in AI.
@@ 3 Evaluation Paradigms
Dimension Benchmarking Evals Construct-Oriented Exploratory Real-World Impact TEVV Archetype Deng et al. (2009) Ganguli et al. (2022) Guinet et al. (2024) Berglund et al. (2024) Collins et al. (2024) Yang et al. (2023a) Indicators Performance Safety/Fairness - Performance/Behaviour Cost/Fairness/Performance Safety/robustness and reliability Dist. summ. Aggregate Extreme/Aggregate Functional Aggregate/Manual Inspection Aggregate Extreme Subject - System System System System System Measurement Observed Observed Latent Observed Observed Observed Task origin - Design Design Design - Design/Operation Protocol Fixed Interactive - - - - Reference Objective Subjective - - Subjective/Rubric Objective Task Mode - Generation - - Generation - Evaluators Researchers Researchers, Deployers Researchers Researchers - - Motivation Comparison Assurance/Understanding Understanding Understanding Comparison Assurance Disciplines - Security, Bio Cognitive Science - Social Sciences Control Theory Raw Number $72$ $13$ $15$ $18$ $4$ $10$ Percentage $57\%$ $10\%$ $12\%$ $14\%$ $3.2\%$ $7.9\%$
@@ 3.1 Benchmarking Paradigm
Benchmarking offers several inherent advantages. First, the use of a standardised test provides a level playing field for all systems being evaluated, while the use of well-defined, objective metrics reduces the ambiguity and allows easy tracking of performance indicators over time. Benchmarks are often publicly available, increasing the transparency of the evaluation process; this can be further improved by reporting the results at the instance level (Burnell et al., 2023). However, the benchmarking paradigm also has notable limitations. Open sharing of benchmarks can lead to data contamination (Zhou et al., 2023), with test data being used for training. Relatedly, repeated testing on the same benchmark may induce overfitting, optimising for specific tests rather than generalised to broader tasks (Fang et al., 2023). Further, benchmark results are inherently tied to a particular distribution of test items, limiting the generalisability of the measurement (Raji et al., 2021). Further discussions on the limitations of benchmarking can be found in Liao et al. (2021).
@@ 3.2 Evals Paradigm
The Evals Paradigm focuses on system safety, often operationalised as a system’s failure to comply with safety specifications or its tendency to exhibit harmful behaviour during tasks that carry risks (e.g., the extent to which it demonstrates so-called “dangerous capabilities” (Shevlane et al., 2023; Phuong et al., 2024)), although fairness is sometimes considered too. This Paradigm often employs extremes analysis, identifying specific (worst) cases. The aim is providing assurance on a system’s safety or gaining understanding on what causes the unsafe behaviour. A common methodology within the Evals Paradigm is “red-teaming”, an interactive and adversarial process where humans or automated agents attempt to “break” or provoke the system to generate undesirable responses. Red-teaming is commonly employed by model deployers (e.g., OpenAI (OpenAI, 2023), Anthropic (Anthropic, 2023)). Generally, tasks are designed in an adversarial way. Evals are predominantly applied to general-purpose AI systems such as LLMs, for which assessing potential risks is particularly relevant due to their broad (and unpredictable) range of applications. Since red-teaming is often conducted by human evaluators, there is typically no objective reference against which the system’s responses are compared. Instead, evaluation relies on subjective human judgment. Archetypal papers from the Evals paradigm are Ganguli et al. (2022)’s work on red-teaming or Kinniment et al. (2023)’s examination of LLM’s ability to self-replicate.
Evals offer a systematic way to identify flaws in AI systems. The failures uncovered by Evals are concrete and actionable, making them amenable to mitigation through additional training, fine-tuning, or reinforcement learning from human feedback (RLHF) (Christiano et al., 2023). However, a major limitation of Evals is that failing to identify flaws does not indicate that a system is safe for deployment (absence of evidence is not evidence of absence). Many evaluated systems (often LLMs) are poorly understood black boxes and can be highly sensitive to small input variations. The ad-hoc and subjective nature of Evals evaluations risks not rigorously accounting for these subtle changes, limiting their reliability as comprehensive safety assessments.
@@ 3.3 Construct-Oriented Paradigm
The Construct-Oriented Paradigm leverages system responses to quantitatively measure underlying “constructs” that describe the system’s behaviour at an abstract level. These constructs are typically based on existing theories of cognitive traits or capabilities, but they can also be solely inferred from system responses (e.g., using factor analysis). To measure these latent constructs, this paradigm often employs a functional link with the observed system behaviour and designed tasks carefully controlling for confounding factors, possibly adapting tasks from the cognitive sciences literature. This paradigm primarily aims to gain a better understanding of self-contained systems, particularly their performance. The works in this paradigm are mostly authored by researchers with a background in psychology/cognitive science. Archetypal examples are Guinet et al. (2024), which applies psychometric methods to infer the ability of LLMs to solve 8th-grade mathematics tests, and Momennejad et al. (2024), which proposes a protocol for evaluating cognitive capabilities in LLMs, operationalising these capabilities through variations in specific tasks to systematically investigate how these variations influence LLM responses.
@@ 3.5 Real-World Impact Paradigm
While most evaluation paradigms seek to assess specific properties of AI systems, the Real-World Impact (RWI) Paradigm measures the impact of AI systems when deployed in the real world. This paradigm leverages techniques from the social and clinical sciences by running (randomised controlled) trials where, most often, assistance by an AI system to humans is considered as an “intervention” whose effect must be quantified, in terms of a change in aggregate performance on the considered task relative to the system’s cost. The goal is therefore comparing different AI systems, or AI-assisted humans to humans alone. Evaluations are generally carried out interactively or on a fixed dataset and, due to the complexity of real-world tasks, human subjective ratings or rubrics may be employed. A common motivation for these evaluations is the comparison of systems situated in the context of real-world applications requiring generation of novel outputs. An archetypal work from this paradigm is Math Converse (Collins et al., 2024), which investigates the perceived helpfulness of LLMs for mathematics. A second representative work is Si et al. (2025)’s study of LLM’s ability to generate novel research ideas; which observes that human reviewers find LLM-generated ideas to be more novel but less feasible than those crafted by humans.
@@ 3.6 TEVV Paradigm
The Test, Evaluation, Verification, and Validation (TEVV) Paradigm draws on methodologies from formal software verification, with its primary focus on ensuring that AI systems behave in a well-defined and predictable way. TEVV is characterised by focusing on observed extreme values of safety or robustness and reliability measures, with the central goal of assurance—providing bounds or guarantees for a minimum or average level of performance under various conditions. To achieve this, TEVV often explicitly operationalises the constructs it aims to measure, formally defining them to reduce uncertainty of the measurement process. It then employs either designed tasks or operational studies, with a variety of protocols. Works in this paradigm commonly deal with Reinforcement Learning and applied fields such as autonomous driving; Yang et al. (2023a) and Mussot et al. (2024) are archetypal examples of TEVV works in these two fields.
@@ 3.7 Evaluations crossing Paradigm Boundaries
Unsurprisingly, we found several evaluation papers do not fit neatly into a single paradigm, but instead bridge multiple paradigms and combine their methodologies (see also Fig. 1). For example, Perez et al. (2022)’s model-written evaluations straddle the Benchmarking and Evals paradigms, using LLMs to generate a large and varied set of questions, thus creating a standardised benchmark, that aim to provide safety assurances in the style typical of Evals. Similarly, CogBench (Coda-Forno et al., 2024) bridges the Capability-oriented and Exploratory paradigms. Here, the authors build a cognitive phenotype of LLMs using psychological experiments designed to assess different constructs. Simultaneously, they explore a number of hypotheses based on anecdotal evidence, including whether RLHF makes LLMs more human-like and the relationship between model size and tendency to exhibit human-like behaviour. These hybrid approaches demonstrate the flexibility of AI evaluation methodologies.
@@ 5 Challenges and Opportunities
We can integrate our insights with existing discussions on gaps in AI evaluation. For instance, Hutchinson et al. (2022) point out the lack of moral evaluations in AI development raising important questions about data consent, the dignity of data workers, and the social responsibilities of AI developers. These factors were not included in our dimensional analysis due to the widespread lack of reporting on such topics. Similarly, Rauh et al. (2024) identify a “risk coverage gap” with many ethical and social risks currently insufficiently addressed. In our framework, evaluations addressing these risks would likely fall under the Real-World Impact Paradigm, which we found to be the least developed. Therefore, Rauh et al. (2024) and our work jointly highlight how this specific niche is unaddressed. Finally, Huang et al. (2024) points out the lack of verifications with provable guarantees for LLMs, which also surfaced from our analysis of the TEVV paradigm. In general, by drawing attention to gaps in the space, we hope to encourage researchers to develop new evaluation methodologies that better address critical issues.

### v2

@@ 1 Introduction
Figure 1: Two-dimensional projection of the surveyed papers based on our framework’s dimensions. The coordinates were obtained using UMAP projection McInnes et al. (2018) of the Jaccard-distance matrix. Each point represents a surveyed paper, whose colour indicates the paradigms it belongs to. We find clusters of papers corresponding to different paradigms. The interactive plot and the code to reproduce it are available here.
@@
Beyond scientific research and industry practices, under-appreciating the breadth of AI evaluation may impact policy, as various policy initiatives ubiquitously rely on “evaluation”. For instance, the EU AI Act states that developers of a general-purpose AI model with “high impact capabilities evaluated on the basis of appropriate technical tools and methodologies” (European Union, 2024, Art. 51(1)) are subject to additional requirements, such as “perform[ing] model evaluation […] with a view to identifying and mitigating systemic risks” (European Union, 2024, Art. 55(1)); thus, “evaluation” refers, in the same document, both to measurement of capabilities and identification of risks. Ongoing efforts [^1] Such as the Code of Practice for general-purpose AI models, https://digital-strategy.ec.europa.eu/en/policies/ai-code-practice. identifying concrete evaluation tools to satisfy those requirements should consider the full spectrum of AI evaluation practices and understand their terminological and methodological differences, to select the appropriate set of techniques for distinct regulatory purposes.
@@ 1.2 Previous Analyses of AI Evaluation
Several recent surveys have explored AI evaluation from specific perspectives. Many of these are focused on LLMs, overviewing benchmarking (Chang et al., 2024) or red-teaming (Lin et al., 2025), studying how ethical aspects are evaluated (Lyu and Du, 2025) or analysing the literature under a verification and validation perspective (Huang et al., 2024). Other surveys include Ruah et al., Rauh et al. (2024), which focuses on the safety of generative AI systems, and Göllner and Tropmann-Frick (2023), which considers responsible AI principles. In contrast, our work focuses on a wide variety of AI systems and evaluation approaches.
Other work discussed the quality of evaluation practice and instruments (Hernandez-Orallo, 2020; Burden, 2024), for instance touching on reproducibility (Burnell et al., 2023), statistical rigour (Gorman and Bedrick, 2019), validity (Subramonian et al., 2023) and representativeness and fairness (Bergman et al., 2023; Göllner and Tropmann-Frick, 2023). We do not focus on these issues here, as they apply broadly to all AI evaluation tools.
Bieger et al. Bieger et al. (2016) proposed a normative framework for evaluating adaptive general-purpose AI, outlining the purposes of evaluation, the properties to be measured, and the challenges involved. In contrast, we take a descriptive approach to AI evaluation, mapping the landscape by identifying the advantages and limitations of different methodologies without prescribing any particular approach. More recently, Cohn et al. Cohn et al. (2022) defined different “facets” of evaluation instruments. While there is some overlap with our dimensions of analysis, several of their facets refer to validity and consistency, properties that are broadly relevant across all paradigms and are hence excluded from our analysis. While Cohn et al. Cohn et al. (2022) apply their framework to 23 evaluation works, neither Bieger et al. Bieger et al. (2016) nor Cohn et al. Cohn et al. (2022) identify distinct paradigms of evaluation or characterise their defining features. Further, these works do not survey the larger and rapidly evolving AI evaluation landscape shaped by recent advancements in AI.
@@ 3 Evaluation Paradigms
Dimension Benchmarking Evals Construct-Oriented Exploratory Real-World Impact TEVV Archetype (Deng et al., 2009) (Ganguli et al., 2022) (Guinet et al., 2024) (Berglund et al., 2024) (Collins et al., 2023) (Yang et al., 2023a) Indicators Performance Safety/Fairness - Performance/Behaviour Cost/Fairness/Performance Safety/robustness and reliability Dist. summ. Aggregate Extreme/Aggregate Functional Aggregate/Manual Inspection Aggregate Extreme Subject - System System System System System Measurement Observed Observed Latent Observed Observed Observed Task origin - Design Design Design - Design/Operation Protocol Fixed Interactive - - - - Reference Objective Subjective - - Subjective/Rubric Objective Task Mode - Generation - - Generation - Evaluators Researchers Researchers, Deployers Researchers Researchers - - Motivation Comparison Assurance/Understanding Understanding Understanding Comparison Assurance Disciplines - Security, Bio Cognitive Science - Social Sciences Control Theory Raw Number $72$ $13$ $15$ $18$ $4$ $10$ Percentage $57\%$ $10\%$ $12\%$ $14\%$ $3.2\%$ $7.9\%$
@@ 3.1 Benchmarking Paradigm
Benchmarking offers several inherent advantages. First, the use of a standardised test provides a level playing field for all systems being evaluated, while the use of well-defined, objective metrics reduces the ambiguity and allows easy tracking of performance indicators over time. Benchmarks are often publicly available, increasing the transparency of the evaluation process; this can be further improved by reporting the results at the instance level (Burnell et al., 2023). However, the benchmarking paradigm also has notable limitations. Open sharing of benchmarks can lead to data contamination (Zhou et al., 2023), with test data being used for training. Relatedly, repeated testing on the same benchmark may induce overfitting, optimising for specific tests rather than generalised to broader tasks (Fang et al., 2023). Further, benchmark results are inherently tied to a particular distribution of test items, limiting the generalisability of the measurement (Raji et al., 2021). Further discussions on the limitations of benchmarking can be found in Liao et al. Liao et al. (2021).
@@ 3.2 Evals Paradigm
The Evals Paradigm focuses on system safety, often operationalised as a system’s failure to comply with safety specifications or its tendency to exhibit harmful behaviour during tasks that carry risks (e.g., the extent to which it demonstrates so-called “dangerous capabilities” (Shevlane et al., 2023; Phuong et al., 2024)), although fairness is sometimes considered too. This Paradigm often employs extremes analysis, identifying specific (worst) cases. The aim is providing assurance on a system’s safety or gaining understanding on what causes the unsafe behaviour. A common methodology within the Evals Paradigm is “red-teaming”, an interactive and adversarial process where humans or automated agents attempt to “break” or provoke the system to generate undesirable responses. Red-teaming is commonly employed by model deployers (e.g., OpenAI (OpenAI, 2023), Anthropic (Anthropic, 2023)). Generally, tasks are designed in an adversarial way. Evals are predominantly applied to general-purpose AI systems such as LLMs, for which assessing potential risks is particularly relevant due to their broad (and unpredictable) range of applications. Since red-teaming is often conducted by human evaluators, there is typically no objective reference against which the system’s responses are compared. Instead, evaluation relies on subjective human judgment. Archetypal papers from the Evals paradigm are Ganguli et al. Ganguli et al. (2022)’s work on red-teaming or Kinniment et al. Kinniment et al. (2024)’s examination of LLM’s ability to self-replicate.
Evals offer a systematic way to identify flaws in AI systems. The failures uncovered by Evals are concrete and actionable, making them amenable to mitigation through additional training, fine-tuning, or reinforcement learning from human feedback (RLHF) (Christiano et al., 2017). However, a major limitation of Evals is that failing to identify flaws does not indicate that a system is safe for deployment (absence of evidence is not evidence of absence). Many evaluated systems (often LLMs) are poorly understood black boxes and can be highly sensitive to small input variations. The ad-hoc and subjective nature of Evals evaluations risks not rigorously accounting for these subtle changes, limiting their reliability as comprehensive safety assessments.
@@ 3.3 Construct-Oriented Paradigm
The Construct-Oriented Paradigm leverages system responses to quantitatively measure underlying “constructs” that describe the system’s behaviour at an abstract level. These constructs are typically based on existing theories of cognitive traits or capabilities, but they can also be solely inferred from system responses (e.g., using factor analysis). To measure these latent constructs, this paradigm often employs a functional link with the observed system behaviour and designed tasks carefully controlling for confounding factors, possibly adapting tasks from the cognitive sciences literature. This paradigm primarily aims to gain a better understanding of self-contained systems, particularly their performance. The works in this paradigm are mostly authored by researchers with a background in psychology/cognitive science. Archetypal examples are Guinet et al. Guinet et al. (2024), which applies psychometric methods to infer the ability of LLMs to solve 8th-grade mathematics tests, and Momennejad et al. Momennejad et al. (2024), which proposes a protocol for evaluating cognitive capabilities in LLMs, operationalising these capabilities through variations in specific tasks to systematically investigate how these variations influence LLM responses.
@@ 3.5 Real-World Impact Paradigm
While most evaluation paradigms seek to assess specific properties of AI systems, the Real-World Impact (RWI) Paradigm measures the impact of AI systems when deployed in the real world. This paradigm leverages techniques from the social and clinical sciences by running (randomised controlled) trials where, most often, assistance by an AI system to humans is considered as an “intervention” whose effect must be quantified, in terms of a change in aggregate performance on the considered task relative to the system’s cost. The goal is therefore comparing different AI systems, or AI-assisted humans to humans alone. Evaluations are generally carried out interactively or on a fixed dataset and, due to the complexity of real-world tasks, human subjective ratings or rubrics may be employed. A common motivation for these evaluations is the comparison of systems situated in the context of real-world applications requiring generation of novel outputs. An archetypal work from this paradigm is Math Converse (Collins et al., 2023), which investigates the perceived helpfulness of LLMs for mathematics. A second representative work is Si et al. Si et al. (2024)’s study of LLM’s ability to generate novel research ideas; which observes that human reviewers find LLM-generated ideas to be more novel but less feasible than those crafted by humans.
@@ 3.6 TEVV Paradigm
The Test, Evaluation, Verification, and Validation (TEVV) Paradigm draws on methodologies from formal software verification, with its primary focus on ensuring that AI systems behave in a well-defined and predictable way. TEVV is characterised by focusing on observed extreme values of safety or robustness and reliability measures, with the central goal of assurance—providing bounds or guarantees for a minimum or average level of performance under various conditions. To achieve this, TEVV often explicitly operationalises the constructs it aims to measure, formally defining them to reduce uncertainty of the measurement process. It then employs either designed tasks or operational studies, with a variety of protocols. Works in this paradigm commonly deal with Reinforcement Learning and applied fields such as autonomous driving; Yang et al. Yang et al. (2023a) and Mussot et al. Mussot et al. (2024) are archetypal examples of TEVV works in these two fields.
@@ 3.7 Evaluations crossing Paradigm Boundaries
Unsurprisingly, we found several evaluation papers do not fit neatly into a single paradigm, but instead bridge multiple paradigms and combine their methodologies (see also Fig. 1). For example, Perez et al. Perez et al. (2022)’s model-written evaluations straddle the Benchmarking and Evals paradigms, using LLMs to generate a large and varied set of questions, thus creating a standardised benchmark, that aim to provide safety assurances in the style typical of Evals. Similarly, CogBench (Coda-Forno et al., 2024) bridges the Capability-oriented and Exploratory paradigms. Here, the authors build a cognitive phenotype of LLMs using psychological experiments designed to assess different constructs. Simultaneously, they explore a number of hypotheses based on anecdotal evidence, including whether RLHF makes LLMs more human-like and the relationship between model size and tendency to exhibit human-like behaviour. These hybrid approaches demonstrate the flexibility of AI evaluation methodologies.
@@ 5 Challenges and Opportunities
We can integrate our insights with existing discussions on gaps in AI evaluation. For instance, Hutchinson et al. Hutchinson et al. (2022) point out the lack of moral evaluations in AI development raising important questions about data consent, the dignity of data workers, and the social responsibilities of developers. These factors were not included in our dimensional analysis due to the widespread lack of reporting on such topics. Similarly, Rauh et al. Rauh et al. (2024) identify a “risk coverage gap” with many ethical and social risks currently insufficiently addressed. In our framework, evaluations addressing these risks would likely fall under the Real-World Impact Paradigm, which we found to be the least developed. Therefore, Rauh et al. Rauh et al. (2024) and our work both highlight how this niche is unaddressed. Finally, Huang et al. Huang et al. (2024) points out the lack of verifications with provable guarantees for LLMs, which also surfaced from our analysis of the TEVV paradigm. In general, by drawing attention to gaps in the space, we hope to encourage researchers to develop new evaluation methodologies that better address issues.
@@ 7 Conclusion
Acknowledgements
JB received funding from the Effective Ventures Foundation Long Term Future Fund through Grant ID: a3rAJ000000017iYA. LP received funding from US DARPA (grant HR00112120007, RECoG-AI) and Open Philanthropy. JHO acknowledges the grants SFERA (PID2021-122830OB-C42) and Cátedra ENIA-UPV, TSI-100930-2023-9 funded by Spain’s MCIN.